OpenStudy (anonymous):

who can solve this Copper; Q= 1kgX387X(30.74-100)= Aluminum; Q= 1kgX899X(30.82-0)=

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

what defines a soltuoin?

OpenStudy (amistre64):

why -100 and -0?

OpenStudy (anonymous):

Mass A intitial temp 100°C thermal equilibrium temp 30.74°C mass 1 kg made of copper Mass B initial temp 0°C thermal temp 30.82°C mass 1 kg made of aluminum

OpenStudy (anonymous):

@amistre64 Q= energy transfer M= mass of substance C= specific heat capacity Delta T= change in temperature Copper; Q= 1kgX387X(30.74-100)= Aluminum; Q= 1kgX899X(30.82-0)=

OpenStudy (anonymous):

that's the answer to your why -100 and -0?

OpenStudy (amistre64):

1kg 899 (30.82 - d) <-- starts out at 30.82, and loses heat 1kg 387 (30.74 + d) <-- start out as 30.74 and gains heat 899(30.82 - d) = 387(30.74 +d) solve for d to determine what the equilibrium temp needs to be. 100 and 0 make no sense at the start

OpenStudy (amistre64):

unless im just not reading the post correctly ...

OpenStudy (amistre64):

we start at 100 and 0? why do you have 4 different temps floating about?

OpenStudy (amistre64):

if we start out as 100 and 0, then they tend to equilize about 30.09, but youve posted different setups and it shard to make any sense of it